steam expands through a well insulated turbine from inlet condition 300c and 4Mpa with negligible velocity to exit condition 40m/s and 0.075Mpa. The turbine is operating at steady state and has an exit diameter of 0.6m. select 2 different values of exit quality in the range of 0.2 to 0.5 and subsequently plot the power generated by the turbine in kW
Answer #1

Steam T2=2 Pa = 0.075 mpa. Ti=300c (= 40 mis D2 = 0.6m. BI = 4 Mpa 0.2 to our a=0 V2=2 P from steam table mama = m = Sanava Am2 = (hf 2 + x (hg-hf 2 - x = 0.3 ha = 384.18 + 0.3 [P278.73 1067.79 kJ/kg 7 * = 0.3 (hfa) + Chile -) * = 0.4 ha 0.4 [ 2278-747kW power Makw ཁཧི་པང་ a royness fraction is 6.3 din 0.5 But power J wear anda 2 소 tear Blades decreases. of